OpenStudy (anonymous):

PLEASE HELP! What number must multiply each side of the equation 2/3x = 16 to produce the equivalent equation x = 24?

OpenStudy (anonymous):

@jackoo

OpenStudy (anonymous):

please help!!

OpenStudy (anonymous):

Divided both of the equations ! 2/3x/x=?

OpenStudy (anonymous):

Or 16/24=? we have same answers in both !

OpenStudy (anonymous):

1.5?

OpenStudy (anonymous):

Nope !:) 16/24=?

OpenStudy (anonymous):

but its gotta be one of these A. 3/2 B. 2/3 C. -2/3 D. -3/2

OpenStudy (anonymous):

its 2/3!!!!!!!!!!

OpenStudy (anonymous):

It B caz : 2/3x/x=2/3 and : 16/24=2/3 :) Got it ?:)

OpenStudy (anonymous):

right?????

OpenStudy (anonymous):

Yep !

OpenStudy (anonymous):

omg thankyou!!!!

OpenStudy (anonymous):

can you help me with one more? plzzzzzz

OpenStudy (anonymous):

il tag you in another question so u can get another medal

OpenStudy (anonymous):

I dont answer for medal !